OpenStudy (anonymous):

In trapezoid ABCD, if the base DC = 16 and the base AB = 20 and E is the midpoint of AD and F is the midpoint of BC, what is the measure of EF?

OpenStudy (amistre64):

regardless of the height of the trap; the midpoint will always be halfway between the a length of 16 and 20; 16+20 ------ = 8+10 = 18 2